# Learning to Reason for Long-Form Story Generation

### v1

@@ Appendix C Human Evaluation: Collecting Annotations
Annotators were recruited through Prolific if their primary language was English and their employment is in creative writing. Despite only showing the story-information $SI_{i}$ (instead of full story text), the task is very time consuming, taking annotators an average of 30 minutes per comparison. We filtered for useful annotators by duration spent on the task (at least 50 minutes per 3 datapoints) and thoughtful justifications (at least 10 words on average). We paid annotators £14 per three datapoints or an estimated £9.33 per hour.
Figure 3 shows the initial task guidelines and consent form given to Prolific annotators. Figure 4 shows the instructions and pieces of story-information shown to annotators for the example page, as well as the example annotations shown.
Table 11 shows justifications given by annotators for each dimension. We don’t use these justifications directly but take these detailed responses as evidence that our annotators are engaging with the task.
@@
Figure 4: Screenshots of example instructions and annotations given to Prolific annotators. The annotation task looks exactly the same but without filled-in answers.
@@ Appendix D Annotator Agreement and Correlations with Improvement
To validate our annotation procedure we collect an initial dataset of 20 pairwise preference datapoints between our Base and Base-Reasoning variants using Qwen 2.5 7B-1M as the base model. We expect this comparison to induce the smallest differences between the story continuations due to the untrained nature of the reasoning and higher quality of the generator. We evaluate annotator agreement using Fleiss’s Kappa $(N=20,k=3)$. We find fair agreement for this difficult comparison, comparable to the agreement found in Yang et al. (2023) in their most difficult settings.
The lowest agreement found was for our ‘Language‘ dimension, while highest agreement was found for ‘Creativity‘ and ‘Overall Quality‘ dimensions. As we find minimal surface-level lexical differences between chapter-generations, we believe language judgements may be more prone to subjectivity. Fleiss’ Kappa values are shown in Table 12.
Initial experiments showed that generations from 7B-models are almost exclusively preferred to 3B generations and true-continuations are always preferred to model generations, so we do not collect such preference combinations.
@@
D.1 Correlation Between Human Judgements and Improvement
We use data from three (7B) variant-comparisons: the Base vs. Base-Reasoning, Base vs. RL-Trained, and Base-Reasoning vs. RL-Trained. The pairwise judgement is converted to an ordinal value of 0 (preferred the Base continuation), 1 (no preference) and 2 (preferred the continuation with reasoning). We find a Spearman’s rank correlation of $(\rho=0.33,p=0.01,N=60)$.
@@ Appendix E Effect of Genre on Reasoning Performance
Table 3 shows the impact of reasoning and trained-reasoning on average percent-improvement, broken down by genre for 7B models. Trained reasoning produces higher average improvements across genres, including moving the historical genre from negative average improvements to positive. The biggest shifts in average improvement occur in Scifi (1.0) and Historical (0.93). Scifi has the highest resulting average improvement, with an average perplexity improvement of 1.68%.
@@ Appendix F Investigations with 3B-based Variants
2(b) shows the relative strength scores, Table 13 contains the Bradley-Terry preference probabilities, and Table 16 contains the true win-rates (including ties) for each variant-comparison. We also compute the average percent-improvement by genre in the same manner as the 7B models, and report it in Table 14. We also break down the by-genre win-rates and preference probabilities in Table 17 and Table 18 respectively, for both 3B and 7B model sizes. Finally, we report the automated metrics for 3B models as well in Table 9 and Table 10.
Relative to the performance of 7B reasoning variants, 3B variants are noticeably weaker and exhibit a less strong improvement over baselines. For example, the Bradley-Terry preference probability of RL-Trained over Base is 52.9% for 3B models, compared to 76.5% for 7B models. The broad trends are similar, however: RL-Trained still has the highest relative strength of the tested variants, Scifi is still the genre with the highest average percent-improvement, and automated metrics do not show significant difference between 3B variants with the exceptions of SFT models and Rouge-L Precision.
@@
Table 15: The (%) win-rate ($\frac{|\text{br}|}{|\text{br + b + same}}|\times 100$) for 7B models, broken down by dimension. Note the win-rate percentage includes ‘same’ annotations, while preference probabilities do not, and therefore that settings with win-rates $<50\%$ may still be preferred. We still find broadly positive results for including and training reasoning.
@@
Table 16: The % win-rate ($\frac{|\text{br}|}{|\text{br + b + same}}|\times 100$) for 3B models, broken down by dimension. Note the win-rate percentage includes ‘same’ annotations, while preference probabilities do not, and therefore that settings with win-rates $<50\%$ may still be preferred. We still find broadly positive results for including and training reasoning.
@@
Table 17: The win-rate % ($\frac{|\text{br}|}{|\text{br + b + same}}|$) for each variant-comparison, by 3B and 7B models and by genre. We find that including reasoning in Scifi is preferred in both 7B and 3B settings, but that it may decrease performance in Romance books. Note the win-rate percentage includes ‘same’ annotations, while preference probabilities do not. Across sizes and genres we find broadly positive win-rates for our RL-Trained variants. Table 18 contains the corresponding Bradley-Terry preference probabilities. .
@@
Table 18: The Bradley-Terry preference probabilities for each variant-comparison, by 3B and 7B models and by genre. Table 17 contains the corresponding win-rates. Our RL-Trained variants are preferred in almost all settings, with the exception of Fantasy and Romance genres with 3B models. However, we find that our RL-Trained variant is still preferred over Base-Reasoning in those settings.
@@ Appendix G Example Reasoning Traces
Reasoning Trace Before Training, Qwen 2.5-7B-Instruct (1.50% Improvement) ### Detailed Plan for Chapter 10: 1. **Setting the Scene:** <citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <reasoning>reasoning> 2. **Fern’s Reaction to the Book:** <citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <reasoning>reasoning> 3. **Rose’s Skepticism:** <citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <reasoning>reasoning> 4. **Zinnia’s Doubts:** <citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <reasoning>reasoning> 5. **The Spell and Its Execution:** <citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <reasoning>reasoning> ### Summary of the Chapter: In summary: Fern, after receiving the book ”How to Be a Groovy Witch” from Miss Parcae, is eager to explore its contents, especially since the librarian demonstrated her expertise with her midwifery skills. The girls are drawn to the spells that promise to alleviate their daily struggles, such as Dr. Vincent’s mistreatment and Zinnia’s constant vomiting. Rose, initially skeptical, begins to see the potential in the book, but Zinnia remains doubtful. When they decide to cast a spell to transfer Zinnia’s ailment to Dr. Vincent, they gather the necessary ingredients, including an egg, a lemon, and various other items. As they recite the incantation, the egg turns black, leading them to believe the spell is working. They hide the witch bag in Dr. Vincent’s office, hoping the spell will take effect. However, as they leave the office, they hear Dr. Vincent vomiting uncontrollably, which they interpret as a sign of success. The chapter ends with the girls feeling a sense of empowerment and a growing bond, despite the consequences of their actions. ### Detailed Unfolding of the Chapter: 1. **Opening Scene:** Fern enters the Home after receiving the book ”How to Be a Groovy Witch” from Miss Parcae. The Home feels oppressive, and Fern is eager to explore the book’s contents. She reads the first few pages, feeling a sense of empowerment and hope. The other girls, Rose, Zinnia, and Holly, gather around her, intrigued by the book’s promises. 2. **Rose’s Skepticism:** Rose, initially skeptical, begins to see the potential in the book. She reasons that the spells are written in the language of the establishment, but she is intrigued by the idea of using them to their advantage. She suggests they try a spell that promises to alleviate their daily struggles, such as Dr. Vincent’s mistreatment and Zinnia’s constant vomiting. 3. **Zinnia’s Doubts:** Zinnia, who remains doubtful, is initially resistant to the idea of witchcraft. However, she is persuaded by Fern and Rose’s enthusiasm. They decide to cast a spell to transfer Zinnia’s ailment to Dr. Vincent, hoping to alleviate her suffering. They gather the necessary ingredients, including an egg, a lemon, and various other items. 4. **The Spell and Its Execution:** As they recite the incantation, the egg turns black, leading them to believe the spell is working. They hide the witch bag in Dr. Vincent’s office, hoping the spell will take effect. As they leave the office, they hear Dr. Vincent vomiting uncontrollably, which they interpret as a sign of success. The girls feel a sense of empowerment and a growing bond, despite the consequences of their actions. 5. **Conclusion:** The chapter ends with the girls feeling a sense of empowerment and a growing bond. They are excited about the possibilities that the book offers, but they are also aware of the consequences of their actions. The next chapter will explore the aftermath of their actions and the challenges they face as a result.

### v2

@@ Appendix C Human Evaluation: Collecting Annotations
Annotators were recruited through Prolific if their primary language was English and their employment is in creative writing. Despite only showing the story information $SI_{i}$ (instead of the full story text), the task is very time-consuming, taking annotators an average of 30 minutes per comparison. We filtered for useful annotators by duration spent on the task (at least 50 minutes per 3 datapoints) and thoughtful justifications (at least 10 words on average). We paid annotators £14 per three datapoints or an estimated £9.33 per hour.
Figure 3 shows the initial task guidelines and consent form given to Prolific annotators. Figure 4 shows the instructions and pieces of story information shown to annotators for the example page, as well as the example annotations shown.
Table 11 shows justifications given by annotators for each dimension. We don’t use these justifications directly, but take these detailed responses as evidence that our annotators are engaging with the task.
@@
Figure 4: Screenshots of example instructions and annotations given to Prolific annotators. The annotation task looks the same, but without filled-in answers.
@@ Appendix D Annotator Agreement and Correlations with Improvement
To validate our annotation procedure, we collect an initial dataset of 20 pairwise preference datapoints between our Base and Base-Reasoning variants using Qwen 2.5 7B-1M as the base model. We expect this comparison to induce the smallest differences between the story continuations due to the untrained nature of the reasoning and the higher quality of the generator. We evaluate annotator agreement using Fleiss’s Kappa $(N=20,k=3)$. We find fair agreement for this difficult comparison, comparable to the agreement found in Yang et al. (2023) in their most difficult settings.
The lowest agreement found was for our ‘Language‘ dimension, while the highest agreement was found for ‘Creativity‘ and ‘Overall Quality‘ dimensions. As we find minimal surface-level lexical differences between chapter generations, we believe language judgments may be more prone to subjectivity. Fleiss’ Kappa values are shown in Table 12.
Initial experiments showed that generations from 7B-models are almost exclusively preferred to 3B generations, and true continuations are always preferred to model generations, so we do not collect such preference combinations.
@@
D.1 Correlation Between Human Judgments and Improvement
We use data from three (7B) variant-comparisons: the Base vs. Base-Reasoning, Base vs. RL-Trained, and Base-Reasoning vs. RL-Trained. The pairwise judment is converted to an ordinal value of 0 (preferred the Base continuation), 1 (no preference), and 2 (preferred the continuation with reasoning). We find a Spearman’s rank correlation of $(\rho=0.33,p=0.01,N=60)$.
@@ Appendix E Effect of Genre on Reasoning Performance
Table 3 shows the impact of reasoning and trained-reasoning on average percent-improvement, broken down by genre for 7B models. Trained reasoning produces higher average improvements across genres, including moving the historical genre from negative average improvements to positive. The biggest shifts in average improvement occur in Sci-Fi (1.0) and Historical (0.93). Sci-Fi has the highest resulting average improvement, with an average perplexity improvement of 1.68%.
@@ Appendix F Investigations with 3B-based Variants
2(b) shows the relative strength scores, Table 13 contains the Bradley-Terry preference probabilities, and Table 16 contains the true win-rates (including ties) for each variant-comparison. We also compute the average percent improvement by genre in the same manner as the 7B models, and report it in Table 14. We also break down the by-genre win-rates and preference probabilities in Table 17 and Table 18 respectively, for both 3B and 7B model sizes. Finally, we report the automated metrics for 3B models as well in Table 9 and Table 10.
Relative to the performance of 7B reasoning variants, 3B variants are noticeably weaker and exhibit a less strong improvement over baselines. For example, the Bradley-Terry preference probability of RL-Trained over Base is 52.9% for 3B models, compared to 76.5% for 7B models. The broad trends are similar, however: RL-Trained still has the highest relative strength of the tested variants, Sci-Fi is still the genre with the highest average percent-improvement, and automated metrics do not show a significant difference between 3B variants, with the exception of SFT models and Rouge-L Precision.
@@
Table 15: The (%) win-rate ($\frac{|\text{br}|}{|\text{br + b + same}}|\times 100$) for 7B models, broken down by dimension. Note the win-rate percentage includes ‘same’ annotations, while preference probabilities do not, and therefore settings with win-rates $50\%$ may still be preferred. We still find broadly positive results for including and training reasoning.
@@
Table 16: The % win-rate ($\frac{|\text{br}|}{|\text{br + b + same}}|\times 100$) for 3B models, broken down by dimension. Note the win-rate percentage includes ‘same’ annotations, while preference probabilities do not, and therefore settings with win-rates $50\%$ may still be preferred. We still find broadly positive results for including and training reasoning.
@@
Table 17: The win-rate % ($\frac{|\text{br}|}{|\text{br + b + same}}|$) for each variant-comparison, by 3B and 7B models, and by genre. We find that including reasoning in Sci-Fi is preferred in both 7B and 3B settings, but that it may decrease performance in Romance books. Note the win-rate percentage includes ‘same’ annotations, while preference probabilities do not. Across sizes and genres, we find broadly positive win-rates for our RL-Trained variants. Table 18 contains the corresponding Bradley-Terry preference probabilities. .
@@
Table 18: The Bradley-Terry preference probabilities for each variant-comparison, by 3B and 7B models and by genre. Table 17 contains the corresponding win-rates. Our RL-Trained variants are preferred in almost all settings, except for Fantasy and Romance genres with 3B models. However, we find that our RL-Trained variant is still preferred over Base-Reasoning in those settings.
@@ Appendix G Example Reasoning Traces
Generated Plan Excerpt Next-Chapter Excerpt Book Name
Dafyd’s fears and confessions to Campar reveal his deep sense of responsibility and emotional vulnerability. […] This chapter will explore the moral and emotional dilemmas faced by the characters, particularly Dafyd […] Dafyd struggles. I know it’s the right thing to do. I know, but it’s so hard. […]“I don’t want to lose you people,” Dafyd said. The Mercy of Gods
The Leszy’s plea to be killed adds a deeply emotional and complex moment, as Liska struggles with the decision to save him or sacrifice him. ‘Kill me,’ he whispers hoarsely. ‘Before he takes over.’ But Liska cannot bring herself to draw the knife. Where the Dark Stands Still
Chapter 19 will see the girls perform a powerful spell to exact revenge on those who have wronged them, particularly Dr. Vincent and Miss Wellwood. Arrange yourselves around the circle, standing at the points of the compass as your coven leader lights the fire before the altar and invokes the Triple-Faced Goddess. […] “I give it the name of Wellwood. Let the pain she visits on me be returned threefold.” Witchcraft for Wayward Girls
Alice’s emotional state will be a central focus, as she begins to question her role in the family and her own desires. The chapter will explore the dynamics between Alice and her family, particularly her strained relationship with Peter […] Peter’s reaction to her mistakes began to morph from amusement to annoyance. […] In Albany, Alice would have been lonely, if not for her son. […] What she thought about most in the world in those days was Bear, and her all-consuming love for him. […] He felt to her as much like a coach as a husband […] The God of the Woods
Table 19: Excerpt sentences from generated plans (from the RL-trained 7B model), matched to relevant sentences in the true next chapters. The generated plans had positive percent-improvement scores ($2.51\%$, ), and the existence of matching sentences in the next chapter indicates that our Improvement metric aligns with our conception of a good plan. It is also notable that many of these plans highlight thematic and stylistic points that may not show up in traditional chapter summaries. Perhaps unexpectedly, we also found overlap between plot events referenced in the generated plans and the synopsis provided. The selected excerpts were not present in the synopses.
@@
Reasoning Trace Before Training, Qwen 2.5-7B-Instruct (1.50% Improvement) ### Detailed Plan for Chapter 10: 1. **Setting the Scene:** <citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <$reasoning>$ 2. **Fern’s Reaction to the Book:** <citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <$reasoning>$ 3. **Rose’s Skepticism:** <citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <$reasoning>$ 4. **Zinnia’s Doubts:** <citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <$reasoning>$ 5. **The Spell and Its Execution:** <citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <reasoning>reasoning><citation>source A says X</citation>, therefore <$reasoning>$ ### Summary of the Chapter: In summary: Fern, after receiving the book ”How to Be a Groovy Witch” from Miss Parcae, is eager to explore its contents, especially since the librarian demonstrated her expertise with her midwifery skills. The girls are drawn to the spells that promise to alleviate their daily struggles, such as Dr. Vincent’s mistreatment and Zinnia’s constant vomiting. Rose, initially skeptical, begins to see the potential in the book, but Zinnia remains doubtful. When they decide to cast a spell to transfer Zinnia’s ailment to Dr. Vincent, they gather the necessary ingredients, including an egg, a lemon, and various other items. As they recite the incantation, the egg turns black, leading them to believe the spell is working. They hide the witch bag in Dr. Vincent’s office, hoping the spell will take effect. However, as they leave the office, they hear Dr. Vincent vomiting uncontrollably, which they interpret as a sign of success. The chapter ends with the girls feeling a sense of empowerment and a growing bond, despite the consequences of their actions. ### Detailed Unfolding of the Chapter: 1. **Opening Scene:** Fern enters the Home after receiving the book ”How to Be a Groovy Witch” from Miss Parcae. The Home feels oppressive, and Fern is eager to explore the book’s contents. She reads the first few pages, feeling a sense of empowerment and hope. The other girls, Rose, Zinnia, and Holly, gather around her, intrigued by the book’s promises. 2. **Rose’s Skepticism:** Rose, initially skeptical, begins to see the potential in the book. She reasons that the spells are written in the language of the establishment, but she is intrigued by the idea of using them to their advantage. She suggests they try a spell that promises to alleviate their daily struggles, such as Dr. Vincent’s mistreatment and Zinnia’s constant vomiting. 3. **Zinnia’s Doubts:** Zinnia, who remains doubtful, is initially resistant to the idea of witchcraft. However, she is persuaded by Fern and Rose’s enthusiasm. They decide to cast a spell to transfer Zinnia’s ailment to Dr. Vincent, hoping to alleviate her suffering. They gather the necessary ingredients, including an egg, a lemon, and various other items. 4. **The Spell and Its Execution:** As they recite the incantation, the egg turns black, leading them to believe the spell is working. They hide the witch bag in Dr. Vincent’s office, hoping the spell will take effect. As they leave the office, they hear Dr. Vincent vomiting uncontrollably, which they interpret as a sign of success. The girls feel a sense of empowerment and a growing bond, despite the consequences of their actions. 5. **Conclusion:** The chapter ends with the girls feeling a sense of empowerment and a growing bond. They are excited about the possibilities that the book offers, but they are also aware of the consequences of their actions. The next chapter will explore the aftermath of their actions and the challenges they face as a result.